# On complexity of restricted fragments of Decision DNNF

### v1

@@ Appendix A Proof of Lemma 11
Let $(u,v)$ be the outgoing edge of $u$ labelled by $j$. Let ${\bf b}_{0}=Proj({\bf b},\mathsf{var}(B_{v}))$. By Lemma 20, ${\bf b}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf c}_{0}={\bf b}_{0}\setminus\{(x,i)\}$. Let ${\bf c}_{1}=Proj({\bf c}_{0},\mathsf{var}(B^{\prime}_{v}))$. Let ${\bf h}_{1}={\bf c}_{1}\cup\{(y,j)\}$.
On the one hand ${\bf h}_{1}\subseteq{\bf h}$. On the other hand $\mathsf{var}(B^{\prime}_{v})\subseteq var(B^{\prime})$ simply by construction. We conclude that ${\bf h}_{1}\subseteq{\bf h}_{0}$. By Lemma 20 applied to $B^{\prime}$ and $B^{\prime}_{v}$, we conclude that an arbitrary extension of ${\bf h}_{1}$ to $\mathsf{var}(B^{\prime})$ is an element of $\mathcal{B^{\prime}}$. It follows that ${\bf h}_{0}\in\mathcal{S}(B^{\prime})$.
@@
$\mathcal{S}(B)|_{\bf g}=\mathcal{S}(B^{\prime}_{1})\times\mathcal{S}(B_{2})$ (2)
@@
The right-hand parts of (2) and (3) are the same so the left-hand parts are implying the statement for the considered case.
@@ Appendix B Proof of Lemma 3
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf a}={\bf g}\cup{\bf h}$. Then ${\bf a}\in\mathcal{S}(B)$. This means that there is ${\bf b}\in\mathcal{A}(B)$ such that ${\bf b}\subseteq{\bf a}$.
By construction,$(x,i)\in{\bf b}$ and ${\bf b}_{0}={\bf b}\setminus\{(x,i)\}$ is an element of $\mathcal{A}(B_{v})$. Hence, $(x,i)\in{\bf a}$. Let ${\bf a}_{0}={\bf a}\setminus\{(x,i)\}$. Clearly, $\mathsf{var}({\bf b}_{0})\subseteq\mathsf{var}(B_{v})\subseteq\mathsf{var}({\bf a}_{0})$. Let ${\bf c}_{0}=Proj({\bf a},var(B_{v}))$. We conclude that ${\bf c}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf d}={\bf c}_{0}\setminus{\bf g}_{0}$. It follows that ${\bf d}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}$.
@@
Conversely, let ${\bf h}_{i}\in\mathcal{S}(B_{u_{i}})|_{{\bf g}_{i}}$ for each $i\in\{1,2\}$. Let ${\bf h}={\bf h}_{1}\cup{\bf h}_{2}$ and let ${\bf a}={\bf g}\cup{\bf h}$. We need to demonstrate that ${\bf a}\in\mathcal{S}$. For each $i\in\{1,2\}$, let ${\bf a_{i}}={\bf g}_{i}\cup{\bf h}_{i}$. By assumption, ${\bf a}_{i}\in\mathcal{S}(B_{u_{i}})$ for each $i\in\{1,2\}$ implying existence of ${\bf b}_{1}\subseteq{\bf a}_{1}$ and ${\bf b}_{2}\subseteq{\bf a}_{2}$ such that ${\bf b}_{1}\in\mathcal{A}(B_{u_{1}})$ and ${\bf b}_{2}\in\mathcal{A}(B_{u_{2}})$. Let ${\bf b}={\bf b}_{1}\cup{\bf b}_{2}$. By construction, ${\bf b}\in\mathcal{A}(B)$. As clearly, ${\bf b}\subseteq{\bf a}$, we conclude that ${\bf a}\in\mathcal{S}(B)$. ∎
@@
Let $B$ be a $\wedge_{d}$-obdd obeying an order $\pi$. Let ${\bf g}$ be an assignment to a prefix of $\pi$. Suppose that the source node $u$ of $B$ is a decision node associated with a variable $x$. Let $i\in\{0,1\}$ be such that $(x,i)\in{\bf g}$ Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{(x,i)\}$. Then $B_{v}$ obeys $\pi_{0}=\pi\setminus\{x\}$, ${\bf g}_{0}$ is an assignment to a prefix of $\pi_{0}$ and $B[{\bf g}]$ is obtained from $B_{v}[{\bf g_{0}}]$ by adding the edge $(u,v)$ labelled with $i$
@@
$B_{v}$ obeys $\pi$ but does not contain $x$ hence clearly $B_{v}$ obeys $\pi_{0}$. Clearly, removal of $x$ from a prefix of $\pi$ results in a prefix of $\pi_{0}$. Therefore, ${\bf g}_{0}$ assigns a prefix of $\pi_{0}$.
@@
    Remove the outgoing edge of $u$ labelled with $2-i$.
@@
We now introduce special notation for the variable $X$ as in Lemma 3. In particular, we let $X_{B}({\bf g})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus\bigcup_{w\in L_{B}({\bf g})}\mathsf{var}(B_{w})$.
@@
As $var(B_{v})\setminus\mathsf{var}({\bf g}_{0})\subseteq\mathsf{var}(B)\setminus\mathsf{var}({\bf g})$, we conclude that $X_{B_{v}}({\bf g}_{0})\subseteq X_{B}({\bf g})$. But what are $X_{B}({\bf g})\setminus X_{B_{v}}({\bf g}_{0})$? They are all the variables of $\mathsf{var}(B)\setminus\mathsf{var}(B_{v})$ but $\mathsf{var}({\bf g})$. Hence the corollary follows.
@@
Let $B$ be a $\wedge_{d}$-obdd obeying an order $\pi$. Let ${\bf g}$ be an assignment to a prefix of $\pi$. Suppose that the source $u$ of $B$ is a conjunction node with children $u_{1}$ and $u_{2}$. Let $V=\mathsf{var}(B)$, $V_{1}=\mathsf{var}(B_{u_{1}})$ and $V_{2}=\mathsf{var}(B_{u_{2}})$. Further on, let $\pi_{1}=\pi[V_{1}]$, $\pi_{2}=\pi[V_{2}]$ (where $\pi[V_{i}]$ is the suborder of $\pi$ induced by $V_{i}$), ${\bf g}_{1}=Proj({\bf g},V_{1})$, ${\bf g}_{2}=Proj({\bf g},V_{2})$. Then, for each $i\in\{1.2\}$, $B_{u_{i}}$ is a $\wedge_{d}$-obdd obeying $\pi_{i}$ and ${\bf g}_{i}$ s an assignment over a prefix of $\pi_{i}$. Furthermore, $B[{\bf g}]$ is obtained from $B_{u_{1}}[{\bf g}_{1}]$ and $B_{u_{2}}[{\bf g}_{2}]$ by adding a conjunction node $u$ with children $u_{1}$ and $u_{2}$.
@@
For each $i\in\{1,2\}$, $B_{u_{i}}$ obeys $\pi$. Hence, $B_{u_{i}}$ obeys the suborder of $\pi$ induced by $\mathsf{var}(B_{u_{i}})$. Assume that say $\mathsf{var}({\bf g}_{1})$ do not form a prefix of $\pi_{1}$. Then there is $x\in\mathsf{var}({\bf g}_{1})$ and $y\in\pi_{1}\setminus\mathsf{var}({\bf g}_{1})$ such that $y<_{\pi_{1}}x$. Then $y<_{\pi}x$ and $y\in\pi\setminus\mathsf{var}({\bf g})$ in contradiction to $\mathsf{var}({\bf g})$ being a prefix of $\pi$.
@@
It is clear from the construction of $B[{\bf g}]$ by Lemma 23 that the set of complete decision nodes of $B[{\bf g}]$ is the union of the sets of such nodes of $B_{1}[{\bf g}_{1}]$ and $B_{2}[{\bf g}_{2}]$. Let $w\in L_{B}({\bf g}]$. Assume w.l.o.g. that $w$ is a node of $B_{1}[{\bf g}_{1}]$ as the minimality cannot be destroyed by moving from a graph to a subgraph, we conclude that $w\in L_{B_{1}}({\bf g}_{1})$.
Conversely, let $w\in L_{B_{1}}({\bf g}_{1})\cup L_{B_{2}}({\bf g}_{2})$. We assume w.l.o.g. that $w\in L_{B_{1}}({\bf g}_{1})$ We note that the construction of $B[{\bf g}]$ by Lemma 23 does not add new edges between vertices of $B_{1}[{\bf g}_{1}]$ and a path from a decision node of $B_{2}[{\bf g}_{2}]$ to a decision node of $B_{1}[{\bf g}_{1}]$ is impossible due to the decomposability of $u$. Therefore, we conclude that $w\in L_{B}({\bf g}]$. ∎
@@
It remains to assume that $u$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be the children of $u$. For $i\in\{1,2\}$, let $B_{i}=B_{u_{i}}$, $V_{i}=\mathsf{var}(B)\cap\mathsf{var}(B_{i})$, ${\bf g}_{i}=Proj({\bf g},V_{i})$, $\pi_{i}=\pi[V_{i}]$. By Lemma 21, for each $i\in\{1,2\}$, $B_{i}$ obeys $\pi_{i}$, ${\bf g}_{i}$ is an assignment over a prefix of $\pi_{i}$, and $\mathcal{S}(B_{i})|_{{\bf g}_{i}}\neq\emptyset$. Therefore, we can apply the induction assumption to each $B_{i}[{\bf g_{i}}]$.

### v2

@@ Appendix A Proof of Lemma 12
Let $(u,v)$ be the outgoing edge of $u$ labelled by $j$. Let ${\bf b}_{0}=Proj({\bf b},\mathsf{var}(B_{v}))$. By Lemma 36, ${\bf b}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf c}_{0}={\bf b}_{0}\setminus\{(x,i)\}$. Let ${\bf c}_{1}=Proj({\bf c}_{0},\mathsf{var}(B^{\prime}_{v}))$. By the induction assumption, ${\bf c}_{1}\in\mathcal{S}(B^{\prime}_{v})$.
Let ${\bf h}_{1}={\bf c}_{1}\cup\{(y,j)\}$. On the one hand ${\bf h}_{1}\subseteq{\bf h}$. On the other hand $\mathsf{var}(B^{\prime}_{v})\subseteq\mathsf{var}(B^{\prime})$ simply by construction. We conclude that ${\bf h}_{1}\subseteq{\bf h}_{0}$. By Lemma 36 applied to $B^{\prime}$ and $B^{\prime}_{v}$, we conclude that an arbitrary extension of ${\bf h}_{1}$ to $\mathsf{var}(B^{\prime})$ is an element of $\mathcal{B^{\prime}}$. It follows that ${\bf h}_{0}\in\mathcal{S}(B^{\prime})$.
@@
$\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\mathcal{S}(B^{\prime}_{1})\times\mathcal{S}(B_{2})$ (7)
@@
The claim now follows from combination of (7) and (8).
@@ Appendix B Proof of Lemma 3
Let ${\bf h}\in\mathcal{S}(B)|_{\bf g}$. Let ${\bf a}={\bf g}\cup{\bf h}$. Then ${\bf a}\in\mathcal{S}(B)$.
By construction, $(x,i)\in{\bf a}$. Let ${\bf a}_{0}={\bf a}\setminus\{(x,i)\}$. Clearly, $\mathsf{var}({\bf b}_{0})\subseteq\mathsf{var}(B_{v})\subseteq\mathsf{var}({\bf a}_{0})$. Let ${\bf c}_{0}=Proj({\bf a}_{0},\mathsf{var}(B_{v}))$. We conclude that ${\bf c}_{0}\in\mathcal{S}(B_{v})$. Let ${\bf d}={\bf c}_{0}\setminus{\bf g}_{0}$. It follows that ${\bf d}\in\mathcal{S}(B_{v})|_{{\bf g}_{0}}$.
@@
Conversely, let ${\bf h}_{i}\in\mathcal{S}(B_{u_{i}})|_{{\bf g}_{i}}$ for each $i\in\{1,2\}$. Let ${\bf h}={\bf h}_{1}\cup{\bf h}_{2}$ and let ${\bf a}={\bf g}\cup{\bf h}$. We need to demonstrate that ${\bf a}\in\mathcal{S}(B)$. For each $i\in\{1,2\}$, let ${\bf a_{i}}={\bf g}_{i}\cup{\bf h}_{i}$. By assumption, ${\bf a}_{i}\in\mathcal{S}(B_{u_{i}})$ for each $i\in\{1,2\}$ implying existence of ${\bf b}_{1}\subseteq{\bf a}_{1}$ and ${\bf b}_{2}\subseteq{\bf a}_{2}$ such that ${\bf b}_{1}\in\mathcal{A}(B_{u_{1}})$ and ${\bf b}_{2}\in\mathcal{A}(B_{u_{2}})$. Let ${\bf b}={\bf b}_{1}\cup{\bf b}_{2}$. By construction, ${\bf b}\in\mathcal{A}(B)$. As clearly, ${\bf b}\subseteq{\bf a}$, we conclude that ${\bf a}\in\mathcal{S}(B)$. ∎
@@
Let $B$ be a $\wedge_{d}$-obdd respecting an order $\pi$. Let $\bf g$} be an assignment to a prefix of $\pi$. Suppose that the source node $u$ of $B$ is a decision node associated with a variable $x$. Let $i\in\{0,1\}$ be such that $(x,i)\in{\bf g}$ Let $(u,v)$ be the outgoing edge of $u$ labelled with $i$. Let ${\bf g}_{0}={\bf g}\setminus\{(x,i)\}$. Then $B_{v}$ respects $\pi_{0}=\pi\setminus\{x\}$, ${\bf g}_{0}$ is an assignment to a prefix of $\pi_{0}$ and $B[{\bf g}]$ is obtained from $B_{v}[{\bf g_{0}}]$ by adding the edge $(u,v)$ labelled with $i$
@@
$B_{v}$ obeys $\pi$ but does not contain $x$ hence clearly $B_{v}$ respects $\pi_{0}$. Clearly, removal of $x$ from a prefix of $\pi$ results in a prefix of $\pi_{0}$. Therefore, ${\bf g}_{0}$ assigns a prefix of $\pi_{0}$.
@@
    Remove the outgoing edge of $u$ labelled with 2i2-i.
@@
We now introduce special notation for the set $X$ as in Lemma 3. In particular, we let $X_{B}({\bf g})=(\mathsf{var}(B)\setminus\mathsf{var}(g))\setminus\bigcup_{w\in L_{B}({\bf g})}\mathsf{var}(B_{w})$.
@@
As $\mathsf{var}(B_{v})\setminus\mathsf{var}({\bf g}_{0})\subseteq\mathsf{var}(B)\setminus\mathsf{var}({\bf g})$, we conclude that $X_{B_{v}}({\bf g}_{0})\subseteq X_{B}({\bf g})$. But what are $X_{B}({\bf g})\setminus X_{B_{v}}({\bf g}_{0})$? They are all the variables of $\mathsf{var}(B)\setminus\mathsf{var}(B_{v})$ but $\mathsf{var}({\bf g})$. Hence the corollary follows.
@@
Let $B$ be a $\wedge_{d}$-obdd respecting an order $\pi$. Let $\bf g$} be an assignment to a prefix of $\pi$. Suppose that the source $u$ of $B$ is a conjunction node with children $u_{1}$ and $u_{2}$. Let $V=\mathsf{var}(B)$, $V_{1}=\mathsf{var}(B_{u_{1}})$ and $V_{2}=\mathsf{var}(B_{u_{2}})$. Further on, let $\pi_{1}=\pi[V_{1}]$, $\pi_{2}=\pi[V_{2}]$ (where $\pi[V_{i}]$ is the order of $V_{i}$ with the elements ordered as in $\pi$), ${\bf g}_{1}=Proj({\bf g},V_{1})$, ${\bf g}_{2}=Proj({\bf g},V_{2})$. Then, for each $i\in\{1.2\}$, $B_{u_{i}}$ is a $\wedge_{d}$-obdd respecting $\pi_{i}$ and ${\bf g}_{i}$ s an assignment over a prefix of $\pi_{i}$. Furthermore, $B[{\bf g}]$ is obtained from $B_{u_{1}}[{\bf g}_{1}]$ and $B_{u_{2}}[{\bf g}_{2}]$ by adding a conjunction node $u$ with children $u_{1}$ and $u_{2}$.
@@
For each $i\in\{1,2\}$, $B_{u_{i}}$ respects $\pi$. Hence, $B_{u_{i}}$ obeys the suborder of $\pi$ induced by $\mathsf{var}(B_{u_{i}})$. Assume that say $\mathsf{var}({\bf g}_{1})$ do not form a prefix of $\pi_{1}$. Then there is $x\in\mathsf{var}({\bf g}_{1})$ and $y\in\pi_{1}\setminus\mathsf{var}({\bf g}_{1})$ such that $y<_{\pi_{1}}x$. Then $y<_{\pi}x$ and $y\in\pi\setminus\mathsf{var}({\bf g})$ in contradiction to $\mathsf{var}({\bf g})$ being a prefix of $\pi$.
@@
It is clear from the construction of $B[{\bf g}]$ by Lemma 39 that the set of complete decision nodes of $B[{\bf g}]$ is the union of the sets of such nodes of $B_{1}[{\bf g}_{1}]$ and $B_{2}[{\bf g}_{2}]$. Let $w\in L_{B}({\bf g})$. Assume w.l.o.g. that $w$ is a node of $B_{1}[{\bf g}_{1}]$ as the minimality cannot be destroyed by moving from a graph to a subgraph, we conclude that $w\in L_{B_{1}}({\bf g}_{1})$.
Conversely, let $w\in L_{B_{1}}({\bf g}_{1})\cup L_{B_{2}}({\bf g}_{2})$. We assume w.l.o.g. that $w\in L_{B_{1}}({\bf g}_{1})$ We note that the construction of $B[{\bf g}]$ by Lemma 39 does not add new edges between vertices of $B_{1}[{\bf g}_{1}]$ and a path from a decision node of $B_{2}[{\bf g}_{2}]$ to a decision node of $B_{1}[{\bf g}_{1}]$ is impossible due to the decomposability of $u$. Therefore, we conclude that $w\in L_{B}({\bf g})$. ∎
@@
It remains to assume that $u$ is a conjunction node. Let $u_{1}$ and $u_{2}$ be the children of $u$. For $i\in\{1,2\}$, let $B_{i}=B_{u_{i}}$, $V_{i}=\mathsf{var}(B)\cap\mathsf{var}(B_{i})$, ${\bf g}_{i}=Proj({\bf g},V_{i})$, $\pi_{i}=\pi[V_{i}]$. By Lemma 37, for each $i\in\{1,2\}$, $B_{i}$ respects $\pi_{i}$, ${\bf g}_{i}$ is an assignment over a prefix of $\pi_{i}$, and $\mathcal{S}(B_{i})|_{{\bf g}_{i}}\neq\emptyset$. Therefore, we can apply the induction assumption to each $B_{i}[{\bf g_{i}}]$.